Question 4 Tries remaining: 2 government imposes a a price ceiling of $68 Points out of 7.70 Calculate the dollar amount of consumer surplus from the price ceiling. The demand for wheat is given by: Q- 186-0.4P. The supply of wheat is given by: Qs- 3P -120. Suppose the Flag question (Do not include a $ sign in your response. Round to the nearest two decimal places if necessary.) Answer: Check
Answer #1

Given, QD = 186-0.4P

And Qs 3P-120

First of all determine the equilibrium price and quantity by equating quantity demanded with quantity supplied

186-0.4P-3P-120

Solving for P we get

186 + 120 = 3P + 0.4P

3.4P = 306

306 3.4

PE = $ 90 per unit

QE = 150 units (=186-0.4*90)

When government imposes a price ceiling of $68

Then quantity demanded will be

QC = 186 -0.4*68 = 158.8 units

When price is 68 Quantity Supplied will be

QSC = 3*68 - 120 = 204 - 120 = 84 UNITS

At a output of 84 units the quantity demanded will be

QD = 186-0.4P

84- 186- 0.4P

0.4P-186-84-102

P = $ 255 per unit

Consumer Surplus = (1/2)*(465-255)*84 +(255-68)*84

= 8820 + 15708

= $ 24,528

Refer the attached picture the consumer surplus = A+B+C

Ap 255 d 31 150 152.